You have purchased 64 shares of Games Inc. for $ 24 and 400 shares of Thrones Co for 670. Over the following three quarters, the companies earned the returns below Quarter 1 Quarter 2 Quarter 3 Games Inc. 20% -23% 35% Thrones Co. -20% 46% -18% What was the buy-and-hold return on your portfolio? Provide your answer in percent, rounded to two decimals, omitting the % sign.
Answer #1

Value of Stock today = 24*64 + 67*400

Value of Stock today = $269,536

Value per Share of Games Inc after 3 quarters = 24(1.20)(0.77)(1.35)

Value per Share of Games Inc after 3 quarters = $29.94

Value of Thrones Co. after 3 quarters = 670(0.80)(1.46)(0.82)

Value of Thrones Co. after 3 quarters = $641.70

Value of Stock after 3 quarters = 29.94*64 + 641.70*400

Value of Stock after 3 quarters = $258,596.16

Return Rate = (258,596.16 - 269,536)/269,536

Return Rate = -4.06%